What is the momentum of a 5000 kg train car that is traveling at 15 m/s?


1
Expert's answer
2021-06-07T09:29:19-0400

By definition, the momentum is the product of the mass and velocity:


"p= mv"

Substituting "m = 5000kg" and "v = 15m\/s", obtain:


"p = 5000kg\cdot 15m/s = 75000\space kg\cdot m/s"

Answer. "75000\space kg\cdot m/s".